OpenStudy (anonymous):

a line passes through the point (2,-4) and has a slope of -4

OpenStudy (anonymous):

y=-4x+4

OpenStudy (triciaal):

y = mx + b using the point (2, -4) find b -4 = -4(2) + b -4 = -8 + b -4 +8 =b b = 4 so as above y = -4x+4